OpenStudy (anonymous):

Find the derivative of y y=sin^-1 (3/t^2)

OpenStudy (freckles):

i suggest finding the derivative implicitly if you don't remember derivative for inverse sine function

OpenStudy (anonymous):

What I got up to so far: y'= -6t^-3/ cos y I dont know where to proceed from there. :/

OpenStudy (freckles):

\[\sin(y)=\frac{3}{t^2} \\ \text{ or } \sin(y)=3t^{-2}\]

OpenStudy (freckles):

you are done unless you want to write y' in terms of t

OpenStudy (freckles):

only

OpenStudy (freckles):

you can do this by making a reference right triangle for the equation sin(y)=3/t^2

OpenStudy (freckles):

 find adj and then find cos(y)

OpenStudy (anonymous):

okay, I got cos y= sqr (t^4-9) / t^2, so do i plug that into y' equals thingy?

OpenStudy (freckles):

you replace cos(y) with that yes

OpenStudy (anonymous):

I got the same answer as the one in my textbook after doing that. Thanks! Now I know how to do this lol

OpenStudy (freckles):

cool stuff